# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

### 31 Mar 2026 v1

@@ Abstract
We study the scaling limit of the one-dimensional lattice Ising–Kac–Kawasaki dynamics under conservative Kawasaki exchange rate. For the Kac coarse-grained field $X_{\gamma}$, we derive a martingale formulation with a discrete conservative drift and a Dynkin martingale. The nonlinear drift is identified by a conservative multiscale replacement scheme based on one-block/two-block estimates, yielding a cubic conservative term in the macroscopic equation. For the stochastic part, we compute the predictable quadratic variation, and obtain a divergence-form Gaussian noise. As a consequence, $X_{\gamma}$ converges to a one-dimensional stochastic Cahn–Hilliard equation with conserved noise.
@@ 1 Introduction
Phase separation with a conserved order parameter is classically described at the continuum level by the Cahn–Hilliard equation, or equivalently by Model B in the Hohenberg–Halperin classification [1, 13, 19]. Its stochastic counterpart is expected to govern mesoscopic fluctuations around phase-separating profiles under a conservation law, and therefore naturally takes the form of a conservative SPDE, namely a stochastic Cahn–Hilliard equation driven by divergence-type noise. In its deterministic form, it can be interpreted as an $H^{-1}$-gradient flow of a Ginzburg–Landau-type free energy, which naturally enforces mass conservation. Random perturbations—motivated by thermal fluctuations and mesoscopic coarse-graining—lead to the stochastic Cahn–Hilliard equation, typically driven by a divergence-form noise so that the conservation law is preserved at the stochastic level [8]. A central problem in nonequilibrium statistical mechanics is to derive such a macroscopic conservative SPDE directly from a microscopic conservative interacting particle system [21].
In this work we consider the one-dimensional Ising–Kac model [14] under Kawasaki exchange dynamics on $\Lambda_{N}:=\mathbb{Z}/(2N+1)\mathbb{Z}$. Let $\Sigma_{N}:=\{-1,1\}^{\Lambda_{N}}.$ We consider a kernel $\kappa_{\gamma}(z)$ of the form with a Kac parameter $\gamma\in(0,\frac{1}{4})$,
@@
The Ising–Kac–Kawasaki dynamics $\sigma(t)$ is a continuous-time Markov chain on $\Sigma_{N}$ generated by nearest-neighbour exchanges: for each bond $(i,i+1)$ let $\sigma^{i,i+1}$ denote the configuration obtained by swapping $\sigma_{i}$ and $\sigma_{i+1}$, and define the heat-bath exchange rate
$c_{i}(\sigma):=\frac{1}{1+\exp\!\bigl(\beta\,\Delta_{i}H_{\gamma}(\sigma)\bigr)},\qquad\Delta_{i}H_{\gamma}(\sigma):=H_{\gamma}(\sigma^{i,i+1})-H_{\gamma}(\sigma).$
The generator is
@@
For a fixed total magnetization $M(N):=\sum_{i\in\Lambda_{N}}\sigma_{i}$, we define the canonical sector
$\Sigma_{N,M}:=\left\{\sigma\in\Sigma_{N}:\ \sum_{i\in\Lambda_{N}}\sigma_{i}=M(N)\right\}.$
@@
where $Z_{N,\gamma,\beta}^{M}:=\sum_{\sigma\in\Sigma_{N,M}}$ is the normalizing constant.
@@
$\sum_{i\in\Lambda_{N}}\sigma_{i}(t)=\sum_{i\in\Lambda_{N}}\sigma_{i}(0)=M(N)\qquad\text{for all }t\geq 0.$
@@
$h_{\gamma}(\sigma;i):=\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\sigma_{j},\quad i\in\Lambda_{N},$
and define the the Kac coarse-grained field $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}h_{\gamma}\!\left(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\right),$ where $\varepsilon_{\gamma}i\in\mathbb{T}$, and $(\alpha_{\gamma},\varepsilon_{\gamma},\delta_{\gamma})$ are the time, space and amplitude scales. Since the Kac kernel is normalized by $\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(z)=1,$ we have $\sum_{i\in\Lambda_{N}}h_{\gamma}(\sigma;i)=\sum_{j\in\Lambda_{N}}\sigma_{j}=M$. Thus for the Kac coarse-grained field $X_{\gamma}$ also has conservation law:
@@
$\langle X_{\gamma}(t),\phi\rangle_{\gamma}=\langle X_{\gamma}(0),\phi\rangle_{\gamma}+\int_{0}^{t}\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma_{s/\alpha_{\gamma}})\,ds+M_{t}^{\gamma}(\phi),$ (1)
@@
$\partial_{t}X=-\nu\Delta^{2}X-A\Delta X+\chi\Delta(X^{3})+\sigma_{*}\nabla\cdot\xi,\quad\int_{\mathbb{T}}X(\cdot,x)dx\equiv M.$ (2)
@@
On the interacting-particle side, hydrodynamic and fluctuation theories for conservative systems have developed a powerful multiscale toolbox based on local equilibrium, the Boltzmann–Gibbs principle, spectral-gap estimates, and one-block/two-block replacements [21, 15, 6]. These methods explain how microscopic conservative exchanges generate macroscopic conservation laws, and they strongly suggest that the right route to stochastic Model B is through a careful closure of the microscopic current and of the local mobility. Beyond weakly interacting settings, quantitative hydrodynamic limits for conservative dynamics with strong, finite-range interactions have also become accessible [17], highlighting the reach of modern two-scale and functional inequality approaches. However, applying this philosophy to the critical Ising–Kac–Kawasaki regime is nontrivial because both the nonlinear drift and the noise must be extracted from conservative bond observables.
@@
The drift term in (1) can be rewritten in conservative form using the microscopic bond current. A direct computation yields an explicit expression for the energy difference $\Delta_{x}H_{\gamma}$ across a swap, which in turn permits a Taylor expansion of the heat-bath rate. This produces a leading nonlinear current term containing the fast local observable $d^{2}_{x}:=(\sigma_{x}-\sigma_{x+1})^{2}$, which is not itself a function of the coarse-grained field and must be closed by a conservative multiscale replacement argument. This yields a closed macroscopic current of the form $X^{2}\nabla X$, and after taking the outer discrete divergence one obtains the cubic conservative drift $\partial_{x}^{2}(X^{3})$.
Moreover, the quadratic variation of the discrete martingales $M^{\gamma}(\phi)$ involves the local mobility observable $q_{i}:=d^{2}_{i}c_{i}(\sigma)=(\sigma_{i}-\sigma_{i+1})^{2}c_{i}(\sigma),$ and this must be closed by the same conservative replacement mechanism in order to produce a deterministic limiting covariance of divergence type. A martingale central limit theorem then yields convergence to a Gaussian limit with covariance $\int_{\mathbb{T}}|\partial_{x}\phi|^{2}dx$, which identifies the noise as the conservative forcing $\partial_{x}\xi$.
@@
Throughout out this paper, assume that following assumptions hold.
@@
    $\varepsilon_{\gamma}:=\frac{1}{N}\to 0,\qquad\frac{\varepsilon_{\gamma}}{\gamma}\to 0,\qquad\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}\to 0,\qquad\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\to 0,\qquad\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\to 0.$
@@
    $\frac{\mathfrak{m}_{2}}{4}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}}\to\nu>0,\quad\lambda_{\gamma}:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\to\lambda>0,\quad\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\rightarrow\sigma_{*}^{2}>0.$
@@
Assume that the assumption $({\bf A})$-$({\bf C})$ hold. After Fourier extension and letting $\gamma\to 0$, the Kac coarse-grained field $X_{\gamma}$ converges in law to $X$ in $L^{2}(0,T;L^{2}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T}))$, where $X$ is the unique solution of the one-dimensional stochastic Cahn–Hilliard equation
$\partial_{t}X=-\nu\Delta^{2}X-\frac{(1+\beta)\lambda}{2}\Delta X+\frac{\lambda\beta}{6}\Delta(X^{3})+\sigma_{*}\nabla\cdot\xi.$ (3)
Moreover, $X(\cdot)$ has conservation law:
$\langle X(t),1\rangle=\langle X(0),1\rangle=M\qquad\text{for all }t\in[0,T].$
The rest of this paper is organized as follows: Section 2 introduces the Ising–Kac–Kawasaki dynamics, the Dirichlet forms, and basic well-posedness results of one-dimensional stochastic Cahn–Hilliard equation. Section 3 rewrites the drift in current form and performs the Taylor expansion leading to the nonlinear current term. Section 4 establish one-block and two-block estimates and derive the second-order current closure. Section 5 computes the quadratic variation of the Dynkin martingale, proves the mobility replacement, and identifies the conservative Gaussian noise. We establish the discrete $H^{-1}$ energy estimates for $X_{\gamma}$ in Section 6 . Finally, Section 7 combines tightness and identification to conclude the proof of the main theorem.
@@ 2.1 Fourier extension from the lattice to the torus
$\Delta_{\varepsilon_{\gamma}}e_{k}=-\lambda_{\gamma}(k)e_{k},\quad\lambda_{\gamma}(k):=\frac{4}{\varepsilon_{\gamma}^{2}}\sin^{2}(\pi k\varepsilon_{\gamma}),\qquad k\in\mathcal{B}_{\gamma}.$
@@
Let $\lambda(k):=(2\pi k)^{4}.$ Then for every $k\in\mathcal{B}_{\gamma}$ one has
@@
For the consistency estimate, note that
$\sin y=y-\frac{y^{3}}{6}+O(y^{5}),\qquad y\to 0.$
Therefore
@@
$\mathrm{Ext}_{\gamma}(P(A_{\gamma})u)=P(\mathcal{A}_{\gamma})\mathrm{Ext}_{\gamma}u.$
@@
The operator $A_{\gamma}$ is diagonal in the discrete Fourier basis, with symbol $\lambda_{\gamma}(k)$. Since $\mathrm{Ext}_{\gamma}$ preserves discrete Fourier coefficients, the equality $\mathrm{Ext}_{\gamma}(A_{\gamma}u)=\mathcal{A}_{\gamma}\mathrm{Ext}_{\gamma}u$ holds mode. The polynomial functional calculus then follows immediately. ∎
@@
Let $(\varrho_{q})_{q\geq-1}$ be a standard smooth dyadic partition of unity on $\mathbb{Z}$, and let
@@
$\|u\|_{B^{\beta}_{p,q}}^{p}:=\left(\sum_{j\geq-1}2^{\beta qj}\|\Delta_{j}u\|_{L^{p}(\mathbb{T})}^{q}\right)^{1/q}..$
@@
Let $\lim_{\gamma\to 0}\nu_{\gamma}=\nu$, $S_{\gamma}(t):=e^{-t\nu_{\gamma}\Delta_{\gamma}^{2}}$, and $S(t):=e^{-t\nu\Delta^{2}}.$ Under Assumption $({\bf A})$, the following hold.

### 05 Apr 2026 v2

@@ Abstract
This paper investigates the scaling limit of one-dimensional lattice Ising–Kac–Kawasaki dynamics. Starting from a martingale formulation for the Kac coarse-grained field $X_{\gamma}$, we decompose the dynamics into a discrete conservative drift and a Dynkin martingale. The nonlinear drift is analyzed via a conservative multiscale replacement scheme based on one-block and two-block estimates, which yields a cubic conservative term in the macroscopic limit. For the stochastic component, we characterize the predictable quadratic variation to obtain a divergence-form Gaussian noise. By establishing uniform $H^{-1}$ energy estimates, we prove that $X_{\gamma}$ converges to a one-dimensional stochastic Cahn–Hilliard equation with conserved noise. Furthermore, we show that the associated canonical equilibrium measure $\mu_{\gamma}$ converges weakly to the $\phi^{4}_{1}$ measure on the conserved-mass hyperplane.
@@ 1 Introduction
Phase separation with a conserved order parameter is classically described at the continuum level by the Cahn–Hilliard equation, or equivalently by Model B in the Hohenberg–Halperin classification [1, 15, 21]. Its stochastic counterpart is expected to govern mesoscopic fluctuations around phase-separating profiles under a conservation law, and therefore naturally takes the form of a conservative SPDE, namely a stochastic Cahn–Hilliard equation driven by divergence-type noise. In its deterministic form, it can be interpreted as an $H^{-1}$-gradient flow of a Ginzburg–Landau-type free energy, which naturally enforces mass conservation. Random perturbations—motivated by thermal fluctuations and mesoscopic coarse-graining—lead to the stochastic Cahn–Hilliard equation, typically driven by a divergence-form noise so that the conservation law is preserved at the stochastic level [10]. A central problem in nonequilibrium statistical mechanics is to derive such a macroscopic conservative SPDE directly from a microscopic conservative interacting particle system [6, 23].
In this work we consider the one-dimensional Ising–Kac model under Kawasaki exchange [16] dynamics on $\Lambda_{N}:=\mathbb{Z}/(2N+1)\mathbb{Z}$. Let $\Sigma_{N}:=\{-1,1\}^{\Lambda_{N}}.$ We consider a kernel $\kappa_{\gamma}(z)$ of the form with a Kac parameter $\gamma\in(0,\frac{1}{4})$,
@@
For each bond $(i,i+1)$, let $\sigma^{i,i+1}$ denote the configuration obtained from $\sigma$ by exchanging the spins at sites $i$ and $i+1$, namely
$(\sigma^{i,i+1})_{j}:=\begin{cases}\sigma_{i+1},&j=i,\\ \sigma_{i},&j=i+1,\\ \sigma_{j},&j\neq i,i+1.\end{cases}$
We also write $d_{i}(\sigma):=\sigma_{i}-\sigma_{i+1}\in\{0,\pm 2\}.$ Then the exchange energy difference is defined by
$\Delta_{i}H_{\gamma}(\sigma):=H_{\gamma}(\sigma^{i,i+1})-H_{\gamma}(\sigma).$
The Ising–Kac–Kawasaki dynamics $\sigma^{N}(t)$ is a continuous-time Markov chain on $\Sigma_{N}$ generated by the energy difference from nearest-neighbour exchanges $\Delta_{i}H_{\gamma}(\sigma)$. The heat-bath exchange rate is defined as
$c_{i}(\sigma):=\frac{1}{1+\exp\!\bigl(\beta\,\Delta_{i}H_{\gamma}(\sigma)\bigr)}..$
Then the Markov generator is
@@
For a fixed total magnetization $M(N):=\sum_{i\in\Lambda_{N}}\sigma(\cdot,i)$, we define the canonical sector
$\Sigma_{N,M}:=\left\{\sigma\in\Sigma_{N}:\ \sum_{i\in\Lambda_{N}}\sigma(i)=M(N)\right\}.$
@@
where $Z_{N,\gamma,\beta}^{M}:=\sum_{\sigma\in\Sigma_{N,M}}\exp\!\bigl(-\beta H_{\gamma}(\sigma)\bigr)\mathbf{1}_{\Sigma_{N,M}}(\sigma)$ is the normalizing constant.
@@
$\sum_{i\in\Lambda_{N}}\sigma^{N}(t,i)=\sum_{i\in\Lambda_{N}}\sigma^{N}(0,i)=M(N)\qquad\text{for all }t\geq 0.$
@@
$h_{\gamma}(\sigma;i):=\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\sigma^{N}(j),\quad i\in\Lambda_{N},$
and define the Kac coarse-grained field $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}h_{\gamma}\!\left(\sigma^{N}\big(\frac{t}{\alpha_{\gamma}}\big);i\right),$ where the space scale $\varepsilon_{\gamma}:=\frac{1}{2N+1}$ so that $\varepsilon_{\gamma}i\in\mathbb{T}$, and $(\alpha_{\gamma},\delta_{\gamma})$ are the time, and amplitude scales. Since the Kac kernel is normalized by $\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(z)=1,$ we have $\sum_{i\in\Lambda_{N}}h_{\gamma}(\sigma;i)=\sum_{j\in\Lambda_{N}}\sigma^{N}(j)=M$. Thus for the Kac coarse-grained field $X_{\gamma}$ also has conservation law:
@@
$\langle X_{\gamma}(t),\phi\rangle_{\gamma}=\langle X_{\gamma}(0),\phi\rangle_{\gamma}+\int_{0}^{t}\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)\Big(\sigma^{N}\big(\frac{s}{\alpha_{\gamma}}\big)\Big)\,ds+M_{t}^{\gamma}(\phi),$ (1)
@@
$\partial_{t}X=-\nu\Delta^{2}X-A\Delta X+\frac{\chi}{3}\Delta(X^{3})+\sigma_{*}\nabla\cdot\xi,\quad\int_{\mathbb{T}}X(\cdot,x)dx\equiv M.$ (2)
@@
In one space dimension, the equilibrium Gibbs measure of stochastic Cahn–Hilliard equation is the canonical $\phi^{4}_{1}$ measure on the conserved-mass hyperplane $V_{M}:=\{\phi\in L^{2}(\mathbb{T}):\langle\phi,1\rangle=M\}$:
$\mu(d\phi):=\frac{1}{Z^{\prime}}\mathbf{1}_{\phi\in V_{M}}\exp\!\left(-\frac{2}{\sigma_{*}^{2}}\int_{\mathbb{T}}\Bigl(-\frac{A}{2}\phi(x)^{2}+\frac{\chi}{12}\phi(x)^{4}\Bigr)dx\right)N(M,\frac{\sigma_{*}^{2}}{2\nu}(-\Delta)^{-1}),$ (3)
where $N(M,\frac{\sigma_{*}^{2}}{2\nu}(-\Delta)^{-1})$ is the Gaussian measure $L^{2}(\mathbb{T})$ with mean $M$ and covariance $\frac{\sigma_{*}^{2}}{2\nu}(-\Delta)^{-1}$ on $V_{M}$, and $Z^{\prime}$ the normalization constant. Equivalently, in formal terms,
$\mu(d\phi)\propto\exp\!\left(-\frac{2}{\sigma_{*}^{2}}\mathcal{F}(\phi)\right)\mathbf{1}_{\{\langle\phi,1\rangle=M\}}\,\mathcal{D}\phi,$
with free energy
$\mathcal{F}(\phi)=\int_{\mathbb{T}}\left(\frac{1}{2}|\nabla\phi(x)|^{2}-\frac{A}{2}\phi(x)^{2}+\frac{\chi}{12}\phi(x)^{4}\right)\,dx.$
In one dimension, the Wick renormalization is unnecessary, and canonical $\phi^{4}_{1}$ measure $\mu$ is the ergodic and time-reversible measure on $V_{M}\subset L^{2}(\mathbb{T})$, see e.g. [7]. Moreover, it is supported on functions of spatial regularity strictly below $C^{1/2}$, has finite polynomial moments, and is the natural invariant (indeed reversible) Gibbs measure associated with the conservative stochastic Cahn–Hilliard/Model B dynamics.
@@
On the interacting-particle side, hydrodynamic and fluctuation theories for conservative systems have developed a powerful multiscale toolbox based on local equilibrium, the Boltzmann–Gibbs principle, spectral-gap estimates, and one-block/two-block replacements [23, 17, 8]. These methods explain how microscopic conservative exchanges generate macroscopic conservation laws, and they strongly suggest that the right route to stochastic Cahn–Hilliard/Model B equation is through a careful closure of the microscopic current and of the local mobility. Beyond weakly interacting settings, quantitative hydrodynamic limits for conservative dynamics with strong, finite-range interactions have also become accessible [19], highlighting the reach of modern two-scale and functional inequality approaches. However, applying this philosophy to the critical Ising–Kac–Kawasaki regime is nontrivial because both the nonlinear drift and the noise must be extracted from conservative bond observables.
@@
The drift term in (1) can be rewritten in conservative form using the microscopic bond current. A direct computation yields an explicit expression for the energy difference $\Delta_{i}H_{\gamma}$ across a swap, which in turn permits a Taylor expansion of the heat-bath rate. This produces a leading nonlinear current term containing the fast local observable $d^{2}_{i}:=(\sigma_{i}-\sigma_{i+1})^{2}$, which is not itself a function of the coarse-grained field and must be closed by a conservative multiscale replacement argument. This yields a closed macroscopic current of the form $X^{2}\nabla X$, and after taking the outer discrete divergence one obtains the cubic conservative drift $\Delta(X^{3})$.
Moreover, the quadratic variation of the discrete martingales $M^{\gamma}(\phi)$ involves the local mobility observable $q_{i}:=d^{2}_{i}c_{i}(\sigma)=(\sigma_{i}-\sigma_{i+1})^{2}c_{i}(\sigma),$ and this must be closed by the same conservative replacement mechanism in order to produce a deterministic limiting covariance of divergence type. A martingale central limit theorem then yields convergence to a Gaussian limit with covariance $\int_{\mathbb{T}}|\nabla\phi|^{2}dx$, which identifies the noise as the conservative forcing $\nabla\xi$.
@@
Throughout this paper, assume that following assumptions hold.
@@
    $\varepsilon_{\gamma}:=\frac{1}{2N+1}\to 0,\qquad\frac{\alpha_{\gamma}}{\gamma}\to 0,\qquad\frac{\varepsilon_{\gamma}}{\gamma}\to 0,\qquad\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}\to 0,\qquad\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\to 0,.$
@@
    $\frac{\mathfrak{m}_{2}}{4}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}}\to\nu>0,\quad\lambda_{\gamma}:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}=\frac{1}{\alpha_{\gamma}(2N+1)^{2}}\to\lambda>0,\quad\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\rightarrow\sigma_{*}^{2}>0.$
@@
Assume that the assumption $({\bf A})$-$({\bf C})$ hold. After Fourier extension and letting $\gamma\to 0$, the Kac coarse-grained field $X_{\gamma}$ converges in law to $X$ in $L^{2}(0,T;L^{2}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T}))$, where $X\in V_{M}$ is the unique solution of the one-dimensional stochastic Cahn–Hilliard equation
$\partial_{t}X=-\nu\Delta^{2}X-\frac{(1+\beta)\lambda}{2}\Delta X+\frac{\lambda\beta}{6}\Delta(X^{3})+\sigma_{*}\nabla\cdot\xi,\quad\int_{\mathbb{T}}X(\cdot,x)dx\equiv M.$ (4)
Moreover, the Kac coarse-grained field $X_{\gamma}$ has the induced equilibrium measure $\mu_{\gamma}:=(\widetilde{X}_{\gamma})_{\#}\mu_{N,\gamma,\beta}.$, so that $\mu_{\gamma}$ converges weakly to the canonical $\phi^{4}_{1}$ measure $\mu$ on the conserved-mass hyperplane $V_{M}:=\{\phi\in L^{2}(\mathbb{T})$, which is defined by (3) with coefficients $A=\frac{(1+\beta)\lambda}{2}$ and $\chi=\lambda\beta/2$.
Throughout the paper, we use the notation $a\lesssim b$ if there exists a constant $C>0$, independent of the variables under consideration, so that $a\leq C\cdot b$, and we denote $a\simeq b$ if $a\lesssim b$ and $b\lesssim a$. We denote $\Delta=\partial_{x}^{2}$. The Fourier transform on the torus $\mathbb{T}^{d}$ is defined with $\hat{f}(k):=(\mathscr{F}_{\mathbb{T}^{d}}f)(k)=\int_{\mathbb{T}^{d}}e^{-2\pi k\cdot x}f(x)dx$, and the inverse Fourier transform on the torus $\mathbb{T}^{d}$ is given by $(\mathscr{F}^{-1}_{\mathbb{T}^{d}}\hat{f})(x)=\sum_{k\in\mathbb{Z}^{d}}e^{2\pi ik\cdot x}\hat{f}(k)$. For any $u\in\mathcal{S}^{\prime}(\mathbb{T}^{d})$ and $j\geq-1$, we denote $\Delta_{j}u$ as the $j$th Littlewood-Paley blocks of $u$. For Besov spaces and Sobolev space with $\alpha\in\mathbb{R}$, we write $C^{\alpha}:=B^{\alpha}_{\infty,\infty}(\mathbb{T}^{d})$ as the Hölder-Besov space, and $H^{\alpha}:=B^{\alpha}_{2,2}(\mathbb{T}^{d})$ as the Sobolev space.
The rest of this paper is organized as follows: Section 2 introduces the Ising–Kac–Kawasaki dynamics, the Dirichlet forms, and basic well-posedness results of one-dimensional stochastic Cahn–Hilliard equation. Section 3 rewrites the drift in current form and performs the Taylor expansion leading to the nonlinear current term. Section 4 establish one-block and two-block estimates and derive the second-order current closure. Section 5 computes the quadratic variation of the Dynkin martingale, proves the mobility replacement, and identifies the conservative Gaussian noise. We establish the $H^{-1}$ energy estimates for $Y_{\gamma}$ in Section 6. Finally, Section 7 combines tightness and identification to conclude the proof of the main theorem.
@@ 2.1 Fourier extension from the lattice to the torus
$\Delta_{\varepsilon_{\gamma}}e_{k}=-\frac{2}{\varepsilon_{\gamma}}\sin(\pi k\varepsilon_{\gamma})e_{k},\qquad k\in\mathcal{B}_{\gamma}.$
@@
Let $\lambda(k):=(2\pi k)^{4},$ and $\lambda_{\gamma}(k):=\frac{4}{\varepsilon_{\gamma}^{2}}\sin^{2}(\pi k\varepsilon_{\gamma})$. Then for every $k\in\mathcal{B}_{\gamma}$ one has
@@
For the consistency estimate, note that $\sin y=y-\frac{y^{3}}{6}+O(y^{5})$, and $y\to 0$. Therefore
@@
$\mathrm{Ext}_{\gamma}(P(A_{\gamma})u)=P(A_{\gamma})\mathrm{Ext}_{\gamma}u.$
@@
The operator $A_{\gamma}$ is diagonal in the discrete Fourier basis, with symbol $\lambda_{\gamma}(k)$. Since $\mathrm{Ext}_{\gamma}$ preserves discrete Fourier coefficients, the equality $\mathrm{Ext}_{\gamma}(A_{\gamma}u)=A_{\gamma}\mathrm{Ext}_{\gamma}u$ holds modewise. The polynomial functional calculus then follows immediately. ∎
@@
Let $(\varrho_{q})_{q\geq-1}$ be a standard smooth dyadic partition of unity with support $\mathcal{A}_{j}\subset\mathbb{Z}$, and let
@@
$\|u\|_{B^{\beta}_{p,q}}:=\left(\sum_{j\geq-1}2^{\beta qj}\|\Delta_{j}u\|_{L^{p}(\mathbb{T})}^{q}\right)^{1/q}..$
@@
Let $\lim_{\gamma\to 0}\nu_{\gamma}=\nu$, $S_{\gamma}(t):=e^{-t\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}}$, and $S(t):=e^{-t\nu\Delta^{2}}.$ Under Assumption $({\bf A})$, the following hold.